OpenStudy (anonymous):

If f(x)={lnx, 01 f(x) is...

OpenStudy (anonymous):

\[f(x) = \left[\begin{matrix}\ln(x) & 0<x \le1 \\ x^2\ln1 & 1<x \le4\end{matrix}\right]\]